Alices restaurant has a total of 205 seats. Th enumber of seats in the non-smoking section is 73 more than twice the number of seats in the smoking section. How many seats are in each section?
Answer by Maths68(1474)

Alices restaurant has a total of 205 seats. Th enumber of seats in the non-smoking section is 73 more than twice the number of seats in the smoking section. How many seats are in each section?


Number of seats in Smoking Section = x
Th number of seats in the non-smoking section is 73 more than twice the number of seats in the smoking section = 2x+73
Total seats = 205
x+(2x+73)=205
x+2x+73=205
3x+73=205
3x=205-73
3x=132
3x/3=32/3
x=44


Number of seats in Smoking Section = x = 44
Number of seats in the non-smoking section = 2x+73 = 2(44)+73 = 88+73 = 161